Find the slope of the line through the following pairs of points.
(-5, -3) and (-5, 2)
In exercises 51 to 54, find the constant of variation k.
m varies directly with n; m = 144 when n = 8.
The revenue for a sandwich shop is directly proportional
to its advertising budget. When the owner spent $2000 a month on advertising, the
revenue was $120,000. If the revenue is now $180,000, how much is the owner
spending on advertising?
Please help these are all different questions that I plainly do not understand.

Answer by stanbon(75887)

Find the slope of the line through the following pairs of points.
(-5, -3) and (-5, 2)
-----------
slope = change in y / change in x = (2--3)/(-5--5) = 5/0 = undefined
----------------------
Find the constant of variation k.
m varies directly with n; m = 144 when n = 8.
m = kn
144 = k*8
k = 144/8 = 18
--------------------
The revenue for a sandwich shop is directly proportional
to its advertising budget. When the owner spent $2000 a month on advertising, the revenue was $120,000. If the revenue is now $180,000, how much is the owner
spending on advertising?
x/180,000 = 2000/120000
x = 180000(1/60) = $3000 (amount owner is spending on advertising)
==============
Cheers,
Stan H.